OpenStudy (anonymous):

how would i graph this x+4y=8

OpenStudy (anonymous):

If you let x=0 you would get 4y=8 or y=2, thus one point would be (0,2). If you let y=0 you would get x=8, thus a second point would be (8,0). You only need two points to graph a line. So plot those two points and draw a line between them and extend it in both directions.

OpenStudy (anonymous):

See attachment

OpenStudy (katrinakaif):

You can graph this by converting into y = mx + b format..so: x + 4y = 8 4y = -x + 8 y = -x/4 + 8/4 y = -x/4 + 2 Now since you know the y-intercept should be 2 make a point on the y-axis at 2. The slope in this case is -1/4 so you would go down one and right 4..and keep going down 1 right 4 until you create a long enough line =)

OpenStudy (anonymous):

Or what K^2 wrote!

OpenStudy (katrinakaif):

LOL K^2 :P

OpenStudy (anonymous):

who's right

OpenStudy (katrinakaif):

Lol we both are, just explained in different terms =)

OpenStudy (anonymous):

Notice that if you let x=0 on K^2's equation that y still equals 2. There are two basic forms for the equation of a line ax+by=c (Yours in this problem) and y=mx+b the one K^2 converted it into. Either will give the same graph.